OpenStudy (anonymous):

Given lim x->4 (x/10+1)=7/5, illustrate the definition of the limit by finding delta-epsilon when epsilon = 1/15. Now, I don't expect anybody to go through the whole proof but I don't know what to do in this case because I already know what epsilon is equal to. If somebody could help me with this little step it would greatly appreciated.

OpenStudy (anonymous):

start with the definition of the limit being \(\frac{7}{5}\)

OpenStudy (anonymous):

that means given \(\epsilon>0\) we can find a \(\delta)\) so that if \(|x-4|<\delta\) then \[|\frac{x}{10}+1-\frac{7}{4}|<\epsilon\]

OpenStudy (anonymous):

in general, \(\delta\) is an expression written in terms of \(\epsilon\) but in your case \(\epsilon\) is a number, namely \(\frac{1}{15}\) so your \(\delta\) will be a number

OpenStudy (anonymous):

Did you mean \[\frac{ x }{ 10 }+\frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

your expression is \(\frac{x}{10}+1\) right?

OpenStudy (anonymous):

and you want to show that the limit is \(\frac{7}{5}\)

OpenStudy (anonymous):

that means that the difference between the two can be made arbitrarily small, i.e that \[|\frac{x}{10}+1-\frac{7}{5}|\] is small

OpenStudy (anonymous):

Ok, got it.

OpenStudy (anonymous):

now it is some algebra

OpenStudy (anonymous):

The hard part =).

OpenStudy (anonymous):

you need to solve for \(\delta\) to make \[|\frac{x}{10}-\frac{2}{5}|<\frac{1}{15}\]

OpenStudy (anonymous):

you need to solve for \(\delta\) to make \[|\frac{x}{10}-\frac{2}{5}|<\frac{1}{15}\]

OpenStudy (anonymous):

\[|\frac{x-4}{10}|<\frac{1}{15}\] \[\frac{1}{10}|x-4|<\frac{1}{15}\] \[|x-4|<\frac{10}{15}=\frac{2}{3}\]

OpenStudy (anonymous):

The part I am struggling with now is the actual proof.

OpenStudy (anonymous):

sorry site froze up on me. but you can see that the algebra is not so bad

OpenStudy (anonymous):

you mean the proof with \(\epsilon\) instead of \(\frac{1}{15}\) ?

OpenStudy (anonymous):

Kind of in reverse. The proof with 1/15 in place of explicitly stating e.

OpenStudy (anonymous):

This is probly more trouble for you that it is worth and I apologize for making this difficult.

OpenStudy (anonymous):

that one i wrote in detail above

OpenStudy (anonymous):

we already have \(\delta=\frac{3}{2}\) works now run the steps backwards

OpenStudy (anonymous):

if \[|x-4|<\frac{2}{3}\] then \[\frac{1}{10}|x-4|<\frac{1}{15}\] and so \[|\frac{x-4}{10}|<\frac{1}{15}\] \[|\frac{x}{10}-\frac{4}{10}|<\frac{1}{15}\] \[|\frac{x}{10}+1-\frac{7}{5}|<\frac{1}{15}\]